OpenStudy (anonymous):

solve the equation using the zero property. -8n(10n-1)=0

jimthompson5910 (jim_thompson5910):

The zero product property is the idea that if two numbers multiply to zero, then one of those numbers has to be zero

jimthompson5910 (jim_thompson5910):

Algebraically, if A*B = 0 then either A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

So how do we use that to break up -8n(10n-1)=0 ?

OpenStudy (anonymous):

either -80n^2=0 or 8=0 ?

jimthompson5910 (jim_thompson5910):

Think of -8n(10n-1) as -8n*(10n-1)

jimthompson5910 (jim_thompson5910):

basically -8n times (10n-1)

jimthompson5910 (jim_thompson5910):

so the first factor is -8n the second factor is 10n-1

OpenStudy (anonymous):

oh gotcha

jimthompson5910 (jim_thompson5910):

So what two equations do you get?